If three cards are removed at random, one at a time without replacement, find the probability that at least one of the cards is a face card?
Answer by ewatrrr(24785)


Hi
three cards removed at random, one at a time 'without replacement'
Given there are 12 face cards(jack, queen,king in each the 4 suits)
40 of the cards are 'not' face cards
P(at least one of the cards is a face card) = 1 - P(no face cards)
1 - P(no face cards) = 1 - 40/52*39/51*38*50
Note: w/o replaceement:
the number of non-face cards and the number of cards is one less when next card is drawn